OpenStudy (emilyjones284):

Ill give a medal and fan everyone who gives an answer...posted below (:

OpenStudy (emilyjones284):

\[\frac{ 1 }{ 2} +\frac{ 3 }{ 2d }= \frac{ -1 }{ d }\]

OpenStudy (zzr0ck3r):

multiply every term by 2d and tell me what you have

OpenStudy (zzr0ck3r):

d+3=-2 d=-5

OpenStudy (emilyjones284):

wait im confused on the multiplying

OpenStudy (emilyjones284):

what would 1/2 times 2d be?

OpenStudy (zzr0ck3r):

2d(1/2)+2d(3/2d)=2d(-1/d) the two cancel in the first term the 2d cancels in the second and the d cancels in the third d+3=-2

OpenStudy (nurali):

1/2+3/2d=-1/d 1*2d+3*2/2*2d=-1/d 2d+6/4d=-1/d Cross Multiply 2d^2+6d=-4d 2d^2+6d+4d=0 2d^2+10d=0 2d(d+5)=0 2d=0 or d+5=0 d=0 or d=-5

OpenStudy (zzr0ck3r):

of course d cant equal 0

OpenStudy (zzr0ck3r):

there is no reason to bring square terms in simply multiply everything by 2d d+3=-2 cant get any easier

OpenStudy (emilyjones284):

oh ok i got it thank you (: how do you know what the lcd is?

OpenStudy (zzr0ck3r):

you have denominators of 2,d,2d

OpenStudy (zzr0ck3r):

so you find something that all three of those divides 2d is that something

OpenStudy (emilyjones284):

why couldnt it be d?